If you could help me out that would be awesome, let me know, thanksChuck, a single taxpayer earns $58,500 in taxable income and $20,800 in interest from an investment in City of Heflin bonds. (Use the U.S tax rate schedule.) Required: a. How much federal tax will he owe? b. What is his average tax rate? c. What is his effective tax rate? d. What is his current marginal tax rate? Complete this question by entering your answers in the tabs below Req A Req B Req C Req D How much federal tax will he owe? (Do not round intermediate calculations. Round Federal tax to 2 decimal places.) Federal tax

Schedule X-Single If taxable income is over: But not over: $ 9.525 $ 38,700 $ 82,500 $157,500 $200,000 $500,000 The tax is: S 9,525 38,700 $ 82,500 $157.500 $200,000 $500,000 10% of taxable income $952.50 plus 12% of the excess over $9,525 $4,453.50 plus 22% of the excess over $38,700 $14,089.50 plus 24% of the excess over $82,500 $32,089.50 plus 32% of the excess over $157,500 $45,689.50 plus 35% of the excess over $200,000 $150,689.50 plus 37% of the excess over $500,000
Answer #1
a
Federal tax = 4453.50+(58500-38700)*22%= $8809.50
b
Average tax rate = Total tax/Taxable income
Average tax rate = 8809.50/58500= 15.06%
c
Effective tax rate = Total tax/Total income
Effective tax rate = 8809.50/79300= 11.11%
d
Current marginal tax rate is 22% for additional income upto $24000 and additional deductions upto $19800